A charged 50 micro faraday  is connected to a50 mH inductor . What is the angular frequency of free oscillation of the circuit?

Grade:9

1 Answers

Rajat
5 years ago
Angular frequency for the free oscillation of an AC circuit is given by 
w(omega) = 1/√(LC)= 1/(50×10-6×50×10-3)1/2
= 1/(25×10-7)1/2=632.455